In an experiment 1.90g of NH3 reacts with 4.96 of O2



4NH3(g) +5O2(g)=4NO(g)+6H2O(g)



Which is a limiting reactant

1
Expert's answer
2022-03-30T21:15:02-0400

Number of moles of NH3= 1.9/17 = 0.11 moles


Stoichiometric mole = number of moles/ mole in the equation = 0.11/4 = 0.0275




Number of moles of oxygen= 4.96/32 = 0.0155


Stoichiometric mole = 0.155/5 = 0.031




The smaller number is 0.0275, thus NH3 is the limiting reagent